OpenStudy (anonymous):

A food company distributes its tomato soup in two cans of different sizes. For the larger can, the diameter has been increased by 30% , and the height remains the same. By what percentage does the volume of the can increase from the smaller can to the larger can? Round your answer to the nearest percent.

OpenStudy (flying_dutchman):

Small can volume = (π*r^2) height Large can volume = (π*1.3r^2) height you can give r and height all the values you want, and compare them, to make it simple, we can try with 1and 2 SCV = (π*1^2) 2 = 6.2831 LCV = (π*1.3(1)^2) = 10.6185 Then you only have to calculate the percentage of the growth. 10.6185 - 6.2831 = 4.3354 6.2831= 100% 4.3354 = 68.96% Percentage of growth = 69 % ok, you could have guessed it was 69%, my result was 68.96 cause I didn't use all the decimal numbers I should have used.